OpenStudy (anonymous):

how to use separable equations to solve: y'(x)=(-2y(x)+19 )^0.5 and y(-1)=-3 then y(x)=?

OpenStudy (zzr0ck3r):

take y(x) = (-2y(x) + 19)^(1/2) so we have y(-1) = (-2*y(-1) + 19)^(1/2) and we know y(-1) = -3 can you get it from here?

OpenStudy (anonymous):

no sry its y'(x)

OpenStudy (zzr0ck3r):

hmm, something odd... you are saying sqrt of something = -3

OpenStudy (zzr0ck3r):

oh i c, one sec

OpenStudy (zzr0ck3r):

so you are doing implicit differentiation in class?

OpenStudy (anonymous):

y(x)'=(-2y(x)+19)^0.5 and y(-1)=-3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its implicit

OpenStudy (anonymous):

my course is ODE, and we have to use separable equations

OpenStudy (anonymous):

that is y'(x)=g(x)h(y) then integral 1/h(y)= integral g(x)dx+c

OpenStudy (anonymous):

y(x)'=(-2y(x)+19)^0.5 dy/((-2y+19)^0.5) =dx

OpenStudy (anonymous):

and then? I dont not how to integral it , its composite, cause it's not y but y(x)

OpenStudy (anonymous):

....

OpenStudy (anonymous):

http://i.imgur.com/CMgMz.gif

OpenStudy (zzr0ck3r):

sec phone

OpenStudy (anonymous):

(-1/2)*du/(sqrt(u))

OpenStudy (anonymous):

u=-2y+19 go.

OpenStudy (anonymous):

u=-2y(x)+19 actually, so du=-2y'(x), this question is not that simple I think

OpenStudy (anonymous):

cause I already try to calculate by using ur method before, but I didn't get the right answer

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

gl with thinking y(x) isn't y.

OpenStudy (anonymous):

its a implicit fct, y(x) is a fct about x but not y

OpenStudy (anonymous):

you've had a year of calc. ? (T/F)

OpenStudy (anonymous):

then I think y'=g(x)h(y) and h(y)=1,g(x)=(-2y(x)+19 )^0.5

OpenStudy (anonymous):

T

OpenStudy (anonymous):

what did you get for y?

OpenStudy (anonymous):

y=integral (-2y(x)+19 )^0.5 dx +C

OpenStudy (anonymous):

"cause I already try to calculate by using ur method before, but I didn't get the right answer" what did you get?

OpenStudy (anonymous):

y=-1/3(-2y+19)^1.5

OpenStudy (anonymous):

not even close

OpenStudy (anonymous):

i am absolutely wrong, and still confused about y(x) and y, so what do you get for y?

OpenStudy (anonymous):

integrate the expression I gave you; set it equal to x+C; solve for y.

OpenStudy (anonymous):

19/2+(-25/(e^2))(1/2)(e^(-2x))

OpenStudy (anonymous):

is that close?

OpenStudy (anonymous):

what's the integral of u^(-1/2) ?

OpenStudy (anonymous):

thx so much! i finally got the right answer, 3/2-x^2/2+4x, actually at first I got so confused with the f(x), I think its different with y, cause it contains x so cant calculate as y, and finally I found that they r just same, so I got the right answer. Really appreciates a lot !!!!!!